OpenStudy (anonymous):

What is the solution set of |2x + 4| < 8?

myininaya (myininaya):

\[-8<2x+4<8\]

OpenStudy (kirbykirby):

|2x-8| < 8 2x-8 < 8 or 2x-8 < -8 x-4 < 4 or x-4 < -4 x < 8 or x < 0 , hence x = (0, 8) This comes from the definition of absolute value saying that |x| = -x if x<0 or |x| = x if x >= 0

OpenStudy (anonymous):

"or" should be an "and" for intersection

OpenStudy (anonymous):

kirby did you make a typo inside the abs value?

OpenStudy (kirbykirby):

isn't it a union?

OpenStudy (kirbykirby):

I always wrote "or" for these types of questions with absolute value inequalities o_O

OpenStudy (anonymous):

-8<2x+4<8 -12<2x<4 -6<x<2 interval notation (-6,2)

OpenStudy (anonymous):

Let k>0, k a real number, let r be a real number, then |x - r| = k is equivalent to -(x-r)=k union x-r=k |x - r| > k is equivalent to -(x-r)>k union x-r>k |x - r| < k is equivalent to -(x-r)<k intersection x-r<k